# Johan Håstad

### v2

@@ 2.3 Decision trees and partial matchings of the grid
Let us first do the case of even side length and start by scetching the argument. We construct a matching in an onion-like fashion using as many pairs along the perimeter as possible. If two dents are at an even distance then it is possible to perfectly match all nodes between the two dents. This happens iff two adjacent dents are of different colors. If two adjacent dents are of the same color we need to match one node with the node towards the interior creating a new dent in the new square which has side lenght two less than the original square. Let us say this more formally.
@@
Suppose the side length is $S$. Take any dent, and assume for concreteness that it is white. Start matching nodes two and two along the perimeter starting with the node next to this dent. This is straightforward until we hit the next dent. If this dent is black we get a perfect match along the perimeter while if it is white we are forced to create a new white dent. We continue on the other side of this dent and go all around the square. We get a new square of side length $S-2$ and some dents. A dent remains iff it is the same color as the dent preceding it.
@@
It is easy to see that in most situation where we apply Lemma 2.12 there are many possible matchings. It will be convenient for us to think of the matching as given by the boundary conditions. This can be done in many ways and any deterministic procedure to go from the set of dents to the matching works for us.
Lemma 2.2 ensures that in any decision tree with question of the form “To which node is $v$ matched?” there is one possible answer such that the partial matching created by the path so far results in a locally consistent matching. By simply erasing any branch which is not locally consistent we maintain the property of each branch being locally consistent.
Lemma 2.3 ensures a slightly more subtle property. Suppose we are given a decision tree, $T$, of small depth and we have a small partial matching, $\tau$ and we want to explore $T$ given $\tau$. Lemma 2.2 makes sure that as long as the sum of the size of $\tau$ and the depth of the decision tree is small, it is always possible to find a locally consistent partner for node queried by $T$. Lemma 2.3 makes sure that any such answer is also present in $T$ and hence it has not already been pruned away. Let us state this as a lemma
Lemma 2.13
@@ 2.4 $t$-evaluations
    The constant true is represented by a 1-tree and the constant 0 is represented by a 0-tree.
@@
    If $F$ is an axiom of the PHP contradiction then $\varphi(F)$ is a 1-tree.
@@
The key property for $t$-evaluations is the following lemma.
Lemma 2.14
Suppose we have a derivation using the rules of Section 2.2 starting with the axioms of the functional onto PHP on the $n\times n$ grid. Let $\Gamma$ be the set of all sub-formulas of this derivation and suppose there is a $t$-evaluation whose range includes $\Gamma$ where $t\leq n/150$. Then each line in the derivation is mapped to a 1-tree. In particular we do not reach a contradiction.
Proof:  This is the lemma that corresponds to Lemma 6.4 of [10]. It relies on two properties, namely that each axiom is represented by a 1-tree, and that the derivation rules preserve this property. The first property is ensured by the local consistency of any branch in a decision tree. The second property follows from the fact that the derivation rules are sound and we never “get stuck” in a decision tree. By this we mean that it always possible to continue a branch in a decision tree keeping the values locally consistent. This is ensured by Lemma 2.2 and Lemma 2.3.
@@
We need sets of three assignments of size $t$ to extendable in a locally consistent way and $493t\leq n/$ 49, ensures this is possible.
When studying the number of lines in a proof where each line is short, an extension of Lemma 2.14 is needed. This was first done in [15] and we rely on the argument in the full version of [11].
In this situation it is not possible to have a single $t$-evaluation whose range is all the sub-formulas that appear in the proof. Instead we will have separate $t$-evaluations for each line. Furthermore, for each line these $t$-evaluations live at the leaves of a decision tree. As these are very similar to what is called $\ell$-common decision trees (as defined in [9]) we call them $\ell$-common $t$-evaluations.
Definition 2.15
A set of formulas $(F_{i})_{i=1}^{M}$ has an $\ell$-common $t$-evaluation if there is a decision tree of depth $\ell$ with the following properties. Take any leaf of this decision tree and the $\tau$ be the partial assignment defined by this path. At this leaf we have a $t$-evaluation $\varphi_{\tau}$ of the formulas $(F^{i}\lceil_{\tau}))_{i=1}^{M}$.
We need an extension of Lemma 2.14. Suppose we are given a proof of PHP and for each line, $\lambda$, let $\Gamma_{\lambda}$ be the set of all sub-formulas of the formula appearing in this line and let $\varphi_{\tau}^{\lambda}$ be the $t$-evaluation given at the leaf defined by $\tau$ that contains $\Gamma_{\lambda}$ it is domain.
Lemma 2.16
Suppose we have a derivation using the rules of Section 2.2 starting with the axioms of the functional onto PHP on the $n\times n$ grid. Suppose each $\Gamma_{\lambda}$ allows a $\ell$-common $t$-evaluation. Suppose $\ell+t\leq n/150$, then each line of the derivation is mapped to a 1-tree in all leaves of its decision tree.
The proof of this not very complicated but needs some notation. The important property is that the copies of the same formula appearing in multiple lines are mapped in a consistent way. Take a formula $F$ appearing on lines $\lambda_{1}$ and $\lambda_{2}$. Take any leaf in the common decision tree of $\lambda_{1}$ defined by an assignment $\tau_{1}$. Take any assignment $\tau_{1}^{\prime}$ consistent with $\tau_{1}$ in the decision tree $\varphi_{\tau_{1}}^{\lambda_{1}}(F)$ leading to a branch labeled $b$. We claim the following.
Lemma 2.17
For any $F$, $\tau_{1}$, $\tau_{1}^{\prime}$ as described above consider the decision tree $T=\varphi_{\tau_{2}}^{\lambda_{2}}(F)$ where $\tau_{2}$ is locally consistent with $\tau_{1}$ and $\tau_{1}^{\prime}$. Then any path in $T$ defined by an assignment $\tau_{2}^{\prime}$ consistent with $\tau_{1}$, $\tau_{1}^{\prime}$ and $\tau_{2}$ leads to a leaf labeled $b$.
Proof:  We prove this by induction over the complexity of $F$. The lemma is true when $F$ is a single variable as this follows from the definition of a $t$-evaluation.
@@
Finally assume that $F=\lor_{i=1}^{m}F_{i}$ and suppose it violates the lemma getting a tuple $\tau_{1},\tau_{1}^{\prime},\tau_{2}$, and $\tau_{2}^{\prime}$ leading to leaves with different values in the two trees. Assume with without loss of generality that $b=1$, the situation being symmetric between the two lines.
By the the definition of $t$-evaluations there must be an $i$ such that $\varphi_{\tau_{1}}^{\lambda_{1}}(F_{i})\lceil_{\tau_{1}^{\prime}}$ is a 1-tree. Similarly for $\lambda_{2}$, we know that $\varphi_{\tau_{2}}^{\lambda_{2}}(F_{i})\lceil_{\tau_{2}^{\prime}}$ is a 0-tree. As the four-tuple of restrictions is consistent and $F_{i}$ is simpler than $F$ this contradicts the inductive assumption.
@@
Lemma 2.17 tells us that the same formula when considered on different lines and under different $t$-evaluations behaves the same. The proof of Lemma 2.16 is now essentially identical to the proof of Lemma 2.14. In this situation we need triplets of three pairs of assignments to be consistent and hence we use $t+\ell\leq n/150$. We omit the details.
@@ 3 Restrictions
We are given an instance of the PHP on the $n\times n$ grid and a restriction, for a suitable parameter $T$, reduces it to a smaller instance on the $(n/T)\times(n/T)$ grid where $(n/T)$ is an odd number.
We divide the grid in to $(n/T)^{2}$ squares, called super-squares, each with side length $T$. Inside each super-square there are $\Delta$ (for a parameter to be fixed) smaller squares that we from now on call “mini-squares”. We pick one mini-square inside each super-square and let these represent the smaller instance. Each super-square has a color as given by a chess board coloring of the reduced instance. For instance the corner super-squares are all white. Each mini-square has the color of its super-square.
Between each mini-square, $s_{i}$ and any mini-square $s_{j}^{\prime}$ in an adjacent super-square we have $3R$ (for a parameter to be chosen) edge disjoint paths, each of even length. We can match exactly all vertices on this path by matching each node of the path to the appropriate adjacent node. We are also interested in matching each node to the other neighbor on the path and in case we need to include one node in each of the two mini-squares to which it is attached creating a dent in the sense of Lemma 2.12. We think of this as using the path as an augmenting path and hence we sometimes refer to these paths as “augmenting paths”. A matching on such a path is of type 0 if it does not include any node from the attached mini-squares and otherwise it is of type 1.
@@
We set up our restrictions such that it is uniquely determined by the values for the variables $y_{P}$. Outside the paths and and the mini-squares we more or less have a fixed matching. This concludes the high level description of a restriction and let us repeat the argument, now giving all formal details.
@@ 3.1 Details of mini-squares and paths
It is convenient to use the concept of a brick which is a square of size $30\times 30$. We think of routing paths through bricks, but in concrete terms such a path is given by two points of attachment of different colors. It the path is of type 0 these two potential dents are matched inside the brick and if the path is of type 1 these two points are matched to the appropriate neighbor in the neighboring brick. We make sure that the there are at most three dents in each side of the brick and of distance at least 12 to any corner and hence we can use Lemma 2.12 to find a matching of the interior. This matching might not look like a path going through the brick but from an abstract point this is a good way to see it.
@@
All but one mini-square have side length $120\Delta R$. To the exterior we have $4\Delta R$ bricks on each side and the interesting part is the middle $2\Delta R$ bricks. Half of these are used to route the $3\Delta R$ paths to the $\Delta$ mini-squares in the super-square in the given direction. The other $\Delta R$ middle bricks are used to make sure that there is enough space to route the paths through different blocks between the mini-squares. The bricks closer to the corners are only used to get the distance to the corners needed by Lemma 2.12.
We have a special single mini-square in the top left corner super-square [^2] This could be any white super-square and choosing this particular super-square is just to make some fixed choice.. It has side length $1120\Delta R+$ 1 which in particular is an odd number but is otherwise like the other mini-squares. We call this the “designated survivor”.
$T$$T$
@@
In the top row, the $n-120R\Delta-1$ nodes outside the designated survivor are matched in a horizontal matching. Similarly the nodes in the leftmost column outside the designated survivor are matched in a vertical matching. This essentially eliminates one row and one column outside the top left super-square and we assume for convenience that $n\equiv 1$ modulo $30$ and we cover the rest of the grid with bricks.
A super-square is a square of bricks of side length $5\Delta^{2}R$ bricks. We need $150\Delta^{2}R$ to be slightly smaller than $T$. The might need a few extra rows and columns for divisibility reasons but for notational simplicity we ignore this difference and simply assume $T=150\Delta^{2}R$. Starting in the top left corner in each super-square we have $\Delta$ mini-squares along the diagonal. This leaves $\Delta^{2}R$ empty rows of bricks in the bottom of each super-square and $\Delta^{2}R$ empty columns at the right. By our placement of the designated survivor and the elimination of the top row and first column also the top left super-square looks essentially the same as other super-squares. Let us describe how to route horizontal paths.
[…]
For each $j$, we reserve $R$ bricks, all with even index, on the right perimeter of $s_{i}$ and the $k$th brick contains three paths that we route straight right to $c_{ij}^{k}$. Similarly we route paths straight left from the $k$th brick in the $i$th part of $s_{j}^{\prime}$ to the same column. This time using odd index bricks. The path is completed by using the suitable part of $c_{ij}^{k}$ to connect the two pieces.
@@
Lemma 3.1
In each brick outside the mini-squares there is at most one vertical and one horizontal path.
Proof:  Let us look at the area to the right of the diagonal in one super-square and to the left of the diagonal in the super-square to its right. The only vertical parts of paths in this area are from the middle segments of the paths. By definition these are in disjoint columns.
The horizontal parts of paths in the this area are from either segment one or segment three. As the former are in even numbered rows of bricks and the latter in odd numbered rows these are also disjoint.

### v3

@@ 2.3 Decision trees and partial matchings of the grid
Let us first do the case of even side length and start by sketching the argument. We construct a matching in an onion-like fashion using as many pairs along the perimeter as possible. Let us say that two dents are “adjacent” if the perimeter is dent-free between these two nodes. If two adjacent dents are at an even distance then it is possible to perfectly match all nodes between the two dents. This happens iff the two adjacent dents are of different colors. If two adjacent dents are of the same color we need to match one node with the node towards the interior creating a new dent in the new square which has side length two less than the original square. Let us say this more formally.
@@
Suppose the side length is $S$. Take any dent, and assume for concreteness that it is white. Start matching nodes two and two along the perimeter starting with the node next to this dent. This is straightforward until we hit the next dent. If this dent is black we get a perfect match along the perimeter while if it is white we are forced to create a new white dent. We continue on the other side of this dent and go all around the square. We get a new square of side length S2S-2 and some dents. A dent remains iff it is the same color as the dent preceding it.
@@
It is easy to see that in most situation where we apply Lemma 2.5 there are many possible matchings. It is convenient for us to think of the matching as given uniquely by the boundary conditions. This can be done in many ways and any deterministic procedure to go from the set of dents to the matching works for us.
Lemma 2.2 ensures that in any decision tree with questions of the form “To which node is $v$ matched?” there is one possible answer such that the partial matching created by the path so far results in a locally consistent matching. By simply erasing any branch which is not locally consistent we maintain the property of each branch being locally consistent.
Lemma 2.4 ensures a slightly more subtle property. Suppose we are given a decision tree, $T$, of small depth and we have a small partial matching, $\tau$ and we want to explore $T$ given $\tau$. Lemma 2.2 makes sure that as long as the sum of the size of $\tau$ and the depth of the decision tree is small, it is always possible to find a locally consistent partner for any node queried by $T$. Lemma 2.4 makes sure that any such answer is also present in $T$ and hence it has not already been pruned away. Let us state this as a lemma
Lemma 2.8.
@@ 2.4 $t$-evaluations
    The constant true is represented by a $1$-tree and the constant false is represented by a $1$-tree.
@@
    If $F$ is an axiom of the $\operatorname{PHP}$ contradiction then $\varphi(F)$ is a $1$-tree.
@@
Remember that in our situation the nodes of the decision tree ask to which other node a certain node $v$ is matched and we also require the assignment along a path in the decision tree to be locally consistent. In this situation, condition 2 follows from the other conditions, but in general this is not the case. The key property for $t$-evaluations is the following lemma.
Lemma 2.9.
Suppose we have a derivation using the rules of Section 2.2 starting with the axioms of the functional onto $\operatorname{PHP}$ on the $n\times n$ grid. Let $\Gamma$ be the set of all sub-formulas of this derivation and suppose there is a $t$-evaluation whose range includes $\Gamma$ where $t\leq n/150$. Then each line in the derivation is mapped to a $1$-tree. In particular we do not reach a contradiction.
Proof 2.10.
This is the lemma that corresponds to Lemma 6.4 of [10]. It relies on two properties, namely that each axiom is represented by a $1$-tree, and that the derivation rules preserve this property. The first property is ensured by the local consistency of any branch in a decision tree. The second property follows from the fact that the derivation rules are sound and we never “get stuck” in a decision tree. By this we mean that it always possible to continue a branch in a decision tree keeping the values locally consistent. This is ensured by Lemma 2.2 and Lemma 2.4.
@@
We need sets of three assignments of size $t$ to be extendable in a locally consistent way and $3t\leq n/50$ ensures that this is possible.
When studying the number of lines in a proof where each line is short, an extension of Lemma 2.9 is needed. This was first done in [16] and we rely on the argument of [12].
In this situation it is not possible to have a single $t$-evaluation whose range is all the sub-formulas that appear in the proof. Instead, we have separate $t$-evaluations for each line. Furthermore, for each line these $t$-evaluations live at the leaves of a decision tree. As these are very similar to what is called $\ell$-common decision trees (as defined in [11]) we call them $\ell$-common $t$-evaluations.
Definition 2.11.
A set of formulas $(F_{i})_{i=1}^{M}$ has an $\ell$-common $t$-evaluation if there is a decision tree of depth $\ell$ with the following properties. Take any leaf of this decision tree and let $\tau$ be the partial assignment defined by this path. At this leaf we have a $t$-evaluation $\varphi_{\tau}$ of the formulas $(F^{i}\lceil_{\tau}))_{i=1}^{M}$.
We need an extension of Lemma 2.9. Suppose we are given a proof of $\operatorname{PHP}$ and for each line, $\lambda$, let $\Gamma_{\lambda}$ be the set of all sub-formulas of the formula appearing in this line and let $\varphi_{\tau}^{\lambda}$ be the $t$-evaluation given at the leaf defined by $\tau$ that contains $\Gamma_{\lambda}$ it is domain.
Lemma 2.12.
Suppose we have a derivation using the rules of Section 2.2 starting with the axioms of the functional onto $\operatorname{PHP}$ on the $n\times n$ grid. Suppose each $\Gamma_{\lambda}$ allows an $\ell$-common $t$-evaluation. Suppose $\ell+t\leq n/300$, then each line of the derivation is mapped to a $1$-tree in all leaves of its decision tree.
The proof of this is not very complicated but needs some notation. The important property is that the copies of the same formula appearing in multiple lines are mapped in a consistent way. Take a formula $F$ appearing on lines $\lambda_{1}$ and $\lambda_{2}$. Take any leaf in the common decision tree of $\lambda_{1}$ defined by an assignment $\tau_{1}$. Take any assignment $\tau_{1}^{\prime}$ consistent with $\tau_{1}$ in the decision tree $\varphi_{\tau_{1}}^{\lambda_{1}}(F)$ leading to a branch labeled $b$. We claim the following.
Lemma 2.13.
For any $F$, $\tau_{1}$, $\tau_{1}^{\prime}$, and $b$ as described above consider the decision tree $T=\varphi_{\tau_{2}}^{\lambda_{2}}(F)$ where $\tau_{2}$ is locally consistent with $\tau_{1}$ and $\tau_{1}^{\prime}$. Then any path in $T$ defined by an assignment $\tau_{2}^{\prime}$ consistent with $\tau_{1}$, $\tau_{1}^{\prime}$ and $\tau_{2}$ leads to a leaf labeled $b$.
Proof 2.14.
We prove this by induction over the complexity of $F$. The lemma is true when $F$ is a single variable as this follows from the definition of a $t$-evaluation.
@@
Finally, assume that $F=\lor_{i=1}^{m}F_{i}$ and suppose it violates the lemma getting a tuple $\tau_{1},\tau_{1}^{\prime},\tau_{2}$, and $\tau_{2}^{\prime}$ leading to leaves with different values in the two trees. Assume with without loss of generality that $b=1$, the situation being symmetric between the two lines.
By the definition of $t$-evaluations there must be an $i$ such that $\varphi_{\tau_{1}}^{\lambda_{1}}(F_{i})\lceil_{\tau_{1}^{\prime}}$ is a $1$-tree. Similarly, for $\lambda_{2}$, we know that $\varphi_{\tau_{2}}^{\lambda_{2}}(F_{i})\lceil_{\tau_{2}^{\prime}}$ is a 0-tree. As the four-tuple of restrictions is consistent and $F_{i}$ is simpler than $F$ this contradicts the inductive assumption.
@@
Lemma 2.13 tells us that the same formula when considered on different lines and under different $t$-evaluations behaves the same. The proof of Lemma 2.12 is now essentially identical to the proof of Lemma 2.9. In this situation we need triplets of three pairs of assignments to be consistent and hence we use $t+\ell\leq n/300$. We omit the details.
@@ 3 Restrictions
We are given an instance of the $\operatorname{PHP}$ on the $n\times n$ grid and a restriction, for a suitable parameter $T$, reduces it to a smaller instance on the $(n/T)\times(n/T)$ grid where $(n/T)$ is an odd number.
We divide the grid in to $(n/T)^{2}$ squares, called super-squares, each with side length $T$. Inside each super-square there are $\Delta$ (for a parameter to be fixed) smaller squares that we from now on call “mini-squares” (of side length larger than 1, but smaller than the super-squares). We pick one mini-square inside each super-square and let these represent the smaller instance. Each super-square has a color as given by a chess board coloring of the reduced instance. For instance the corner super-squares are all white. Each mini-square has the color of its super-square.
Between each mini-square, $s_{i}$ and any mini-square $s_{j}^{\prime}$ in an adjacent super-square we have $3R$ (for a parameter to be chosen) edge-disjoint paths, each of even length. We can match exactly all vertices on any such path by matching each node of the path to the appropriate adjacent node. We are also interested in matching each node to the other neighbor on the path and in this case we need to include one node in each of the two mini-squares to which it is attached creating a dent in the sense of Lemma 2.5. We think of this as using the path as an augmenting path and hence we sometimes refer to these paths as “augmenting paths”. A matching on such a path is of type 0 if it does not include any node from the attached mini-squares and otherwise it is of type $1$.
@@
Let us point out that for mini-squares in super-squares on the perimeter of the entire grid, in some direction(s) there are no adjacent mini-square. In this situation, no paths attach on the side with no neighbor.
@@
We set up our restrictions such that it is uniquely determined by the values for the variables $y_{P}$. Outside the paths and the mini-squares we more or less have a fixed matching. This concludes the high level description of a restriction and let us repeat the argument, now giving all formal details.
@@ 3.1 Details of mini-squares and paths
It is convenient to use the concept of a brick which is a square of size $30\times 30$. We think of routing paths through bricks, but in concrete terms such a path is given by two points of attachment of different colors. If the path is of type 0 these two potential dents are matched inside the brick and if the path is of type $1$ these two points are matched to the appropriate neighbor in the neighboring brick. We make sure that the there are at most three dents in each side of the brick and of distance at least $12$ to any corner and hence we can use Lemma 2.5 to find a matching of the interior. This matching might not look like a path going through the brick but from an abstract point this is a good way to see it.
@@
Mini-squares.
All but one mini-square have side length $120\Delta R$ (or, equivalently, $4\Delta R$ bricks). We have $4\Delta R$ bricks on each side and the interesting part is the middle $2\Delta R$ bricks. The bricks close to the corners are only used to ensure that we can perfectly match the interior of the mini-square according to Lemma 2.5.
Half of the middle bricks are used to route the $3\Delta R$ paths to the $\Delta$ mini-squares in the super-square in the given direction. The other $\Delta R$ middle bricks are used to make sure that there is enough space to route the paths through different blocks between the mini-squares.
Designated survivor.
We have a special single mini-square in the top left corner super-square [^2] This could be any white super-square and choosing this particular super-square is just to make some fixed choice.. It has side length $120\Delta R+1$ which in particular is an odd number but is otherwise like the other mini-squares. We call this the “designated survivor”.
TTTT
@@
The first row and first column.
In the top row, the $n-120R\Delta-1$ nodes outside the designated survivor are matched in a horizontal matching. Similarly, the nodes in the leftmost column outside the designated survivor are matched in a vertical matching. This essentially eliminates one row and one column outside the top left super-square and we assume for convenience that $n\equiv 1$ modulo $30$ and we cover the rest of the grid with bricks.
Structure of super-squares.
A super-square is a square of bricks of side length $5\Delta^{2}R$ bricks. We need $150\Delta^{2}R$ to be slightly smaller than $T$. We might need a few extra rows and columns for divisibility reasons but for notational simplicity we ignore this difference and simply assume $T=150\Delta^{2}R$. Starting in the top left corner in each super-square we have $\Delta$ mini-squares along the diagonal. This leaves $\Delta^{2}R$ empty rows of bricks in the bottom of each super-square and $\Delta^{2}R$ empty columns at the right. By our placement of the designated survivor and the elimination of the top row and first column also the top left super-square looks essentially the same as other super-squares.
Structure of paths.
Fix a mini-square $s_{i}$ in a super-square $S$ and let us see how to route paths to mini-squares $s_{j}^{\prime}$ in the super-square to its right. For each pair $(i,j)$ we reserve $R$ columns, $c_{ij}^{k},\ 1\leq k\leq R$, of bricks in the right part of $S$. This can be done has we have $\Delta^{2}$ pairs of mini-squares and $\Delta^{2}R$ empty columns of bricks.
We number the middle bricks on each side from one to $2\Delta R$ and for each $j$ we reserve $R$ bricks, all with even number on the right perimeter of $s_{i}$. The $k$th brick contains three paths that we route straight right to $c_{ij}^{k}$. Similarly we route paths straight left from the $k$th brick in the $i$th part of $s_{j}^{\prime}$ to the same column. This time using odd numbered bricks. The path is completed by using the suitable part of $c_{ij}^{k}$ to connect the two pieces.
@@
Lemma 3.1.
In each brick outside the mini-squares there is at most one set of three vertical paths and at most one set of three horizontal paths.
Proof 3.2.
Let us look at the area to the right of the diagonal in one super-square and to the left of the diagonal in the super-square to its right. The only vertical parts of paths in this area are from the middle segments of the paths. By definition these are in disjoint columns.
The horizontal parts of paths in this area are from either segment one or segment three. As the former are in even numbered rows of bricks and the latter in odd numbered rows these are also disjoint.